### v1

# Learning Multiplex Embeddings on Text-rich Networks with One Text Encoder

Abstract
In real-world scenarios, texts in a network are often linked by multiple semantic relations (e.g., papers in an academic network are referenced by other publications, written by the same author, or published in the same venue). Text documents and their relations form a multiplex text-rich network. Mainstream text representation learning methods use pretrained language models (PLMs) to generate one embedding for each text unit, expecting that all types of relations between texts can be captured by these single-view embeddings. However, this presumption does not hold particularly in multiplex text-rich networks. Along another line of work, multiplex graph neural networks (GNNs) directly initialize node attributes as a feature vector for node representation learning, but they cannot fully capture the semantics of the nodes’ associated texts. In this paper, we propose Metern, a new framework for learning Multiplex Embeddings on TExt-Rich Networks. In contrast to existing methods, Metern uses one text encoder to model the shared knowledge across relations and leverages a small number of parameters per relation to derive relation-specific representations. This allows the encoder to effectively capture the multiplex structures in the network while also preserving parameter efficiency. We conduct experiments on nine downstream tasks in five networks from both academic and e-commerce domains, where Metern outperforms baselines significantly and consistently. Code is available at https://github.com/PeterGriffinJin/METERN-submit.
1 Introduction
Texts in the real world are often interconnected by multiple types of semantic relations Jin et al. (2023a). For example, papers connected through “same-venue” relation edges lean towards sharing coarse-grained topics, while papers connected through the “cited-by” relation edges tend to share fine-grained topics; E-commerce items linked by “co-viewed” edges usually have related functions, while items linked by “co-brand” edges can have similar designs. The texts and multiple types of links together form a type of network called a multiplex text-rich network, where documents are treated as nodes and the edges reflect multiplex relations among documents. Given a multiplex text-rich network, it is appealing to learn node representations tailored for different relation types, which can be broadly utilized in various downstream tasks (e.g., paper classification and recommendation in academic networks, and item recommendation and price prediction in e-commerce networks) Jin et al. (2023b, c).
A straightforward way to learn node representations in a text-rich network is by encoding the nodes’ associated texts with pretrained language models (PLMs) Brown et al. (2020); Devlin et al. (2019); Liu et al. (2019b). Many studies Cohan et al. (2020); Ostendorff et al. (2022); Reimers and Gurevych (2019) propose to finetune PLMs with contrastive objectives, pushing the text representations with similar semantics to be close in the latent space while pulling those unrelated apart. The semantics correlation between text units is then measured by the similarity of their representations such as with cosine similarity or dot product. However, most existing approaches based on PLMs use a single vector for each text unit, with the implicit underlying assumption that the semantics of different relations between text units are largely analogous, which does not hold universally, particularly in multiplex text-rich networks (demonstrated in Section 3).
To capture the diverse relationships among nodes, multiplex representation learning Jing et al. (2021); Park et al. (2020); Qu et al. (2017); Zhang et al. (2018) is proposed in the graph domain. The philosophy is to learn multiple representations for every node, each encoding the semantics of one relation. They mainly adopt graph neural networks (GNNs) as the backbone model architecture and utilize separate encoders to encode each relation. Nevertheless, these studies with GNNs represent the texts associated with each node as bag-of-words or context-free embeddings Mikolov et al. (2013), which are not enough to characterize the contextualized text semantics. For example, “llama” in biology books and “llama” in machine learning papers should have different meanings given their context, but they correspond to the identical entry in a bag-of-words vector and possess the same context-free embedding. To capture contextualized semantics, a straightforward idea is to adopt PLMs Brown et al. (2020); Devlin et al. (2019); Liu et al. (2019b). However, it is inefficient and unscalable to have separate PLMs for each relation, given that one PLM usually has millions or billions of parameters Kaplan et al. (2020).
To this end, we propose Metern to learn multiplex node/text embeddings with only one shared PLM encoder. The key idea is to introduce “relation prior tokens”, which serve as priors for learning text embeddings. Specifically, relation prior tokens are added upon the original text tokens and fed into a text encoder. The text encoder is encouraged to learn the shared knowledge across different relations, while the relation prior tokens are propelled to capture relation-specific signals. We further explore how to apply Metern to downstream tasks under different scenarios: (1) direct inference where the source relation is clearly singular (e.g., using the “same-author” relation alone for author identification), and (2) indirect inference where source relations might be hybrid and need to be selected via learning. We evaluate Metern on five large-scale networks from the academic domain and the e-commerce domain with nine downstream tasks, where Metern outperforms competitive baseline methods significantly and consistently.
To summarize, our main contributions are as follows:
  • Conceptually, we identify the semantic shift across different relations and formulate the problem of multiplex representation learning on text-rich networks.
  • Methodologically, we propose Metern, which learns multiplex text representations with one text encoder and multiple relation prior tokens. Then, we introduce direct inference and “learn-to-select-source-relation” inference with Metern on different downstream tasks.
  • Empirically, we conduct experiments on nine downstream tasks on five datasets from different domains, where Metern outperforms competitive baselines significantly and consistently.
Figure 1: Model Framework Overview. Metern has a language model encoder to model the shared knowledge among relations and relation-prior embeddings to capture the relation-specific signals.
2 Preliminaries
2.1 Multiplex Text-rich Networks
In a multiplex text-rich network, each node is associated with texts, and nodes are connected by multiple types of edges. We view the texts in each node as a document, and all such documents constitute a corpus $\mathcal{D}$.
Definition 1
(Multiplex Text-rich Networks) A multiplex text-rich network is defined as $\mathcal{G}=\{\mathcal{G}^{1},\mathcal{G}^{2},...,\mathcal{G}^{|\mathcal{R}|}\}=(\mathcal{V},\mathcal{E},\mathcal{D},\mathcal{R})$, where $\mathcal{G}^{r}=(\mathcal{V},\mathcal{E}^{r},\mathcal{D})$ is a network of the relation type $r\in\mathcal{R}$, $\mathcal{V}$ is the set of nodes, $\mathcal{E}=\bigcup_{r\in\mathcal{R}}\mathcal{E}^{r}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of all edges, $\mathcal{D}$ is the set of documents, and $\mathcal{R}$ is the relation type set. Each node $v_{i}\in\mathcal{V}$ is associated with a document $d_{i}\in\mathcal{D}$. Note that $|\mathcal{R}|>1$ for multiplex networks.
2.2 Problem Definitions
Definition 2
(Learning Multiplex Embeddings on Text-rich Networks) Given a multiplex text-rich network $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{D},\mathcal{R})$, the task of learning multiplex embeddings is to build a model $f_{\Theta}:\mathcal{V}\rightarrow\mathbb{R}^{|\mathcal{R}|\times d}$ with parameters $\Theta$ to learn node representation vectors ${\bm{h}}_{v_{i}}\in\mathbb{R}^{|\mathcal{R}|\times d}$ for each node $v_{i}\in\mathcal{V}$, which should be able to be broadly utilized to various downstream tasks. Note that we aim to learn $|\mathcal{R}|$ embeddings for each node $v_{i}$, with each corresponding to a relation $r\in\mathcal{R}$.
Figure 2: Distribution shift between different relations on Geology network. cb, sa, sv, cr, and ccb represent cited-by, same-author, same-venue, co-reference, and co-cited-by relation respectively. Each entry is the PREC@1 of BERT embedding fine-tuned on the corresponding source relation distribution and tested on the corresponding target relation distribution.
3 Why multiplex embeddings?
Learning a single embedding for each node/document presumes that the embedding ${\bm{h}}_{v}$ is enough to capture the semantics proximity between nodes with a similarity function $\text{Sim}(\cdot)$, i.e., $P(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}},{\bm{h}}_{v_{j}})$, having an underlying assumption that all relations have analogous semantic distributions, i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$ for different relations $r_{k}$ and $r_{l}$. However, this assumption does not always hold in real-world multiplex text-rich networks. For example, two papers ($v_{i},v_{j}$) written by the same author (relation: $r_{k}$) can be either published in the same venue (relation:$r_{l}$) or not; two items ($v_{i},v_{j}$) frequently purchased by the same user (relation: $r_{k}$) can have either the same brand (relation:$r_{l}$) or not.
In addition, we empirically find that the distribution shift across different relations, i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\neq P_{r_{l}}(e_{ij}|v_{i},v_{j})$, in multiplex text-rich networks, truly affects the learned embedding. We finetune BERT [^1] We use the bert-base-uncased checkpoint. Devlin et al. (2019) to generate embeddings on one source relation distribution $P_{r_{k}}(e_{ij}|v_{i},v_{j})$ (row) and test the embeddings on the same or another target relation distribution $P_{r_{l}}(e_{ij}|v_{i},v_{j})$ (column). The results (measured by in-batch PREC@1) are shown in Figure 2. If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly the same, which is not the case in Figure 2. As a result, if we only represent each node/text with one embedding vector, the single representation will mix up the semantics across different relation distributions and will lose accuracy in each of them. This inspires us to learn multiplex embeddings for each node/text, one for each relation, to capture the distinct semantics for each relation. The embeddings belonging to different relations can benefit different downstream tasks, e.g., “same-author” for author identification, and “same-venue” for venue recommendation. More studies on the raw data distributions and learned embeddings can be found in Appendix A.4.
4 Proposed Method
A straightforward way to learn multiplex embeddings is to have multiple encoders Jing et al. (2021); Park et al. (2020), one for each relation. However, this will make the total number of parameters $|\mathcal{R}|$ times as many as that of a single-encoder model, increasing both memory complexity and time complexity for training and inference, especially when every single encoder is computationally expensive (e.g., pretrained language models Devlin et al. (2019); Liu et al. (2019b)). From another perspective, the semantics of different relations can complement each other, e.g., papers citing each other (i.e., connected via the “cited-by” relation) are highly likely to be of similar topics, thus appearing in the same venue (i.e., also connected via the “same-venue” relation). Nevertheless, adopting separate encoders for each relation will prevent the relations from benefiting each other. To this end, we propose to use only one text encoder to learn multiplex embeddings, with a simple but effective design of “relation prior tokens”. Then, we discuss how to apply the learned embeddings to downstream tasks, including direct inference for tasks where the source relation is clearly singular and “learn to select source relations” for tasks where the source relations might be hybrid. The whole model framework can be found in Figure 1.
4.1 Metern: Learning Multiplex Embeddings with One Text Encoder
Representation Learning with Text Encoder.
Mainstream representation learning models Cohan et al. (2020); Reimers and Gurevych (2019) usually adopt a text encoder $\text{Enc}(\cdot)$ (e.g., Transformer Vaswani et al. (2017)) to generate one node/text embedding for node $v_{i}$, i.e., ${\bm{h}}_{v_{i}}=\text{Enc}(d_{i})$. Then, the predicted similarity score between node $v_{i}$ and $v_{j}$ is calculated by
$P(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}},{\bm{h}}_{v_{j}})=\text{Sim}(\text{Enc}(d_{i}),\text{Enc}(d_{j})),$ (1)
where $\text{Sim}(\cdot,\cdot)$ is a similarity measurement function, e.g., dot product Karpukhin et al. (2020) or cosine similarity Reimers and Gurevych (2019).
However, as we have discussed in Section 3, $P_{r_{k}}(e_{ij}|v_{i},v_{j})\neq P_{r_{l}}(e_{ij}|v_{i},v_{j})$ for different relations $r_{k}$ and $r_{l}$. This motivates us to obtain relation-conditioned embeddings ${\bm{h}}_{v|r}$ rather than just ${\bm{h}}_{v}$ to capture the diverse distribution $P_{r}(e)$ conditioned on relation $r$.
Generating Relation-conditioned Embeddings.
We represent the relations by introducing relation prior tokens $P_{r}=\{p^{(1)}_{r},...,p^{(m)}_{r}\}$ with embeddings $Z_{r}=\{{\bm{z}}^{(1)}_{r},...,{\bm{z}}^{(m)}_{r}\}$ for each relation $r$. Each ${\bm{z}}^{(t)}_{r}\in Z_{r}$ is the $k$-dimensional embedding for the relation prior token $p^{(t)}_{r}\in P_{r}$. To generate relation-conditioned embeddings with one encoder, we provide the encoder with both the node information and the target relation, i.e., $d_{i}$ and $P_{r}$, as input. The representation ${\bm{h}}_{v_{i}|r}$ is obtained by
${\bm{h}}_{v_{i}|r}=\text{Enc}(\{P_{r};d_{i}\}),$ (2)
[…]
Then the relation prediction score is calculated by
$P_{r}(e_{ij}|v_{i},v_{j})\ \propto\ \text{Sim}({\bm{h}}_{v_{i}|r},{\bm{h}}_{v_{j}|r})=\text{Sim}(\text{Enc}(\{P_{r};d_{i}\}),\text{Enc}(\{P_{r};d_{j}\})).$ (3)
In our experiment, we adopt the dot product as the similarity calculation function, i.e., $\text{Sim}({\bm{h}}_{v_{i}|r},{\bm{h}}_{v_{j}|r})={\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{j}|r}$.
Multi-Relation Learning Objective.
During the unsupervised representation learning phase, all the relations will be learned simultaneously with the following log-likelihood:
$\max_{\Theta}\mathcal{O}=\sum_{r\in\mathcal{R}}\sum_{e_{ij}\in\mathcal{E}^{r}}\text{log}P_{r}(e_{ij}|v_{i},v_{j};\Theta),$ (4)
Here, the conditional probability $P_{r}(e_{ij}|v_{i},v_{j};\Theta)$ is calculated as follows Oord et al. (2018):
$P_{r}(e_{ij}|v_{i},v_{j};\Theta)=\frac{\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{j}|r})}{\sum_{v_{u}\in\mathcal{V}}\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{u}|r})},$ (5)
To make the calculation efficient, we leverage the negative sampling technique Jin et al. (2020); Mikolov et al. (2013) to simplify the objective and obtain our loss below.
$\min_{\Theta}\mathcal{L}=\sum_{r\in\mathcal{R}}w_{r}\sum_{e_{ij}\in\mathcal{E}^{r}}-\log\frac{\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{j}|r})}{\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{j}|r})+\sum_{v_{u}^{\prime}}\exp({\bm{h}}_{v_{i}|r}\cdot{\bm{h}}_{v_{u}^{\prime}|r})}=\sum_{r\in\mathcal{R}}w_{r}\mathcal{L}_{r}.$ (6)
In the equation above, $v_{u}^{\prime}$ stands for a random negative sample. In our implementation, we use “in-batch negative samples” Karpukhin et al. (2020) to reduce the encoding cost. Note that we add relation weights $w_{r}$ to control the relative learning speed of different relations (analysis can be found in Section 5.8).
4.2 Inference with Metern
Node embeddings of high quality should be able to generalize to various downstream tasks within the given network scenario Hamilton et al. (2017). In this section, we propose to conduct direct inference and “learn-to-select-source-relations” inference with the learned multiplex embeddings.
Direct Inference with an Evident Source Relation.
We propose direct inference for those downstream tasks where the semantically closest source relation $r_{\text{target}}\in\mathcal{R}$ to the target downstream task is clearly singular (e.g.,same-author” relation for author identification and “same-venue” relation for venue recommendation). In these cases, we can directly infer the embedding of $r_{\text{target}}$ (with $P_{r_{\text{target}}}$) for the downstream task, without any downstream task training samples:
${\bm{h}}_{\text{target}}={\bm{h}}_{v|r_{\text{target}}}=\text{Enc}(\{P_{r_{\text{target}}};d\}).$ (7)
Learning to Select Source Relations.
We propose to learn to select source relations with some downstream task training samples when the semantically closest source relation is not clear (e.g., paper classification and year prediction in academic networks; item classification and price prediction in e-commerce networks). We introduce a set of learnable query embeddings for the target task $Q_{\text{target}}=\{{\bm{q}}^{(1)}_{\text{target}},{\bm{q}}^{(2)}_{\text{target}},...,{\bm{q}}^{(s)}_{\text{target}}\}$ to learn to select the source relation embeddings from $Z_{\mathcal{R}}=\{Z_{r}\}_{r\in\mathcal{R}}$ via attention-based mix-up, as follows:
${\bm{z}}^{(t)}_{\text{target}}=\sum_{{\bm{z}}_{i}\in Z_{\mathcal{R}}}\alpha^{(t)}_{i}\cdot{\bm{z}}_{i},\quad\quad\alpha^{(t)}_{i}=\text{softmax}_{{\bm{z}}_{i}\in Z_{\mathcal{R}}}({\bm{z}}_{i}\cdot{\bm{q}}^{(t)}_{\text{target}})$ (8)
$Z_{\text{target}}=\{{\bm{z}}^{(1)}_{\text{target}},{\bm{z}}^{(2)}_{\text{target}},...,{\bm{z}}^{(s)}_{\text{target}}\},\quad\quad P_{\text{target}}=\{p^{(1)}_{\text{target}},p^{(2)}_{\text{target}},...,p^{(s)}_{\text{target}}\},$ (9)
${\bm{h}}_{\text{target}}=\text{Enc}(\{P_{\text{target}};d\}).$ (10)
$Z_{\mathcal{R}}=\{Z_{r}\}_{r\in\mathcal{R}}$ is the source relation embedding pool (with ||×$|\mathcal{R}|\times m$ embeddings), which is the union of all the source relation embeddings for $r\in\mathcal{R}$. We use $P_{\text{target}}$ and $Z_{\text{target}}$ to represent the downstream task prior tokens and their embeddings respectively (${\bm{z}}^{(t)}_{\text{target}}$ is the embedding for $p^{(t)}_{\text{target}}$). Note that only the parameters in $Q_{\text{target}}$ are trainable to learn to select source relations.
With the mechanism above, the model will learn to assign a larger weight $\alpha$ to the relation embedding ${\bm{z}}_{i}$ which can contribute more to solving the downstream task, thus learning to select source relations for downstream tasks. At the same time, the learned weight $\alpha$ can also help reveal the hidden correlation between relations and tasks (See the analysis in Section 5.5 and Appendix A.7).
4.3 Discussions
Complexity Analysis.
In our experiment, we use the Transformer encoder Vaswani et al. (2017) as the $\text{Enc}(\cdot)$ function. Time Complexity: Given a node $v_{i}$ associated with a document $d_{i}$ containing $p$ tokens, the time complexity of our multiplex representation learning model with $m$ embeddings for each relation is $\mathcal{O}((p+m)^{2})$, which is on par with the $\mathcal{O}(p^{2})$ complexity of the single relation representation learning model since $m\ll p$. Memory Complexity: Given a network with $|\mathcal{R}|$ types of relations and $T$ parameters in the $\text{Enc}(\cdot)$, the parameter complexity of our multiplex representation learning model is $\mathcal{O}(T+|\mathcal{R}|mk)$, which is nearly the same with the complexity of $\mathcal{O}(T)$ of the single relation representation learning model since $|\mathcal{R}|mk\ll T$. The empirical time efficiency study and memory efficiency study are shown in Section 5.6.
Difference from Existing Works Lester et al. (2021); Liu et al. (2021); Qin and Eisner (2021).
These works propose to add learnable “prompt tokens” onto the text sequences, before feeding them into the Transformer architecture, sharing a similar design with our relation prior tokens. However, there are three inherent differences between these works and our work: 1) Different focuses: Existing works Lester et al. (2021); Liu et al. (2021); Qin and Eisner (2021) focus more on efficient language model tuning (i.e., training the prompt embedding only), while we focus on joint relation learning (i.e., using the relation embeddings to capture relation specific information and the shared encoder to capture cross-relation shared knowledge). 2) Different designs: The prompt tokens used in NLP problems Lester et al. (2021); Liu et al. (2021); Qin and Eisner (2021) generally correspond to natural language descriptions of the tasks and labels. In our case, we use relation prior tokens to encode the abstract relations between documents that may not be directly described by natural language tokens. As a result, the representation of an unseen relation may need to be learned as a mixture of source relation representations instead of simply being represented by natural language prompts. 3) Different applications: Existing works focus on NLU Liu et al. (2021); Qin and Eisner (2021), NLG Lester et al. (2021), and knowledge probe Qin and Eisner (2021) tasks, while our work focuses on multiplex representation learning.
5 Experiments
In this section, we first introduce the five datasets. Then, we demonstrate the high quality of the learned multiplex embeddings by Metern. After that, we show the effectiveness of Metern on downstream tasks with direct source relation inference and “learn-to-select-source-relation” inference, respectively. Finally, we conduct efficiency analysis, multiplex embedding visualization, and the study of relation weight $w_{r}$. More experiments on relation token embedding initialization and relation token embedding visualization can be found in Appendix A.6 and A.8, respectively.
5.1 Datasets
We run experiments on both academic networks from the Microsoft Academic Graph (MAG) Sinha et al. (2015); Zhang et al. (2023) and e-commerce networks from Amazon He and McAuley (2016). In academic networks, nodes correspond to papers and there are five types of relations among papers: “cited-by” (cb), “same-author” (sa), “same-venue” (sv), “co-reference” (cr), and “co-cited-by” (ccb); while in e-commerce networks, nodes are items and there are four types of relations between items: “co-purchased” (cop), “co-viewed” (cov), “bought-together” (bt), and “co-brand” (cob). Since both MAG and Amazon have multiple domains, we select two domains from MAG and three domains from Amazon. In total, we have five datasets in the experiments (i.e., MAG-Geology, MAG-Mathematics, Amazon-Clothes, Amazon-Home, and Amazon-Sports). The datasets’ statistics can be found in Table 7.
5.2 Baselines
We compare Metern with three kinds of baselines, large-scale corpora finetuned text embedders, multiplex graph neural networks, and multi-relation learning language models. The first category includes SPECTER Cohan et al. (2020), SciNCL Ostendorff et al. (2022), Sentence-Transformer Reimers and Gurevych (2019), and the OpenAI-ada-002 embedder Brown et al. (2020). SPECTER Cohan et al. (2020) is a text embedder finetuned from SciBERT Beltagy et al. (2019) with academic domain contrastive learning. SciNCL Ostendorff et al. (2022) further improves SPECTER by introducing controlled nearest neighbor sampling. For Sentence-Transformer Reimers and Gurevych (2019), we use the recent best checkpoint [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 which finetunes MPNet Song et al. (2020) on above 1 billion sentence pairs from 32 domains. OpenAI-ada-002333https://openai.com/blog/new-and-improved-embedding-model is the recent large language model-based text embedder proposed by OpenAI. The second category includes DMGI Park et al. (2020) and HDMI Jing et al. (2021). DMGI Park et al. (2020) is a graph convolutional network-based multiplex node embedding encoder. HDMI Jing et al. (2021) further improves DMGI by proposing an alternative semantic attention-based fusion module. For multiplex GNN methods, we use bag-of-words embeddings as the initial node feature vectors. The third category includes the vanilla finetuned language model (Vanilla FT) and MTDNN Liu et al. (2019a). For Vanilla FT, we finetuned BERT Devlin et al. (2019) with all the relation pairs, without distinguishing among them. In this way, the model will only output one single embedding for each node. MTDNN Liu et al. (2019a) is a multi-task learning model with a shared language model backbone for different tasks and task-specific modules connected to the backbone. As a result, it can output multiplex representations for each node corresponding to different tasks (relations). Note that Vanilla FT, MTDNN, and Metern are all initialized by the same bert-base-uncased checkpoint.
Table 1: Multiplex representation learning experiments on academic networks: Geology and Mathematics. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively.
Geology Mathematics
Model cb sa sv cr ccb Avg. cb sa sv cr ccb Avg.
SPECTER 12.84 12.89 1.5 5.56 9.1 8.38 28.74 23.55 2.39 15.96 25.59 19.25
SciNCL 15.91 14.3 1.57 6.41 10.4 9.72 36.14 26.41 2.83 19.82 30.69 23.18
MPNet-v2 30.87 20.94 1.94 10.36 17.16 16.25 46.12 29.92 3.11 23.60 36.42 27.83
OpenAI-ada-002 30.39 21.08 2.02 16.57 16.69 17.35 39.86 27.22 2.67 19.81 31.62 24.24
DMGI 28.99 27.79 4.91 9.86 16.32 17.58 46.55 42.62 6.11 27.80 38.87 28.85
HDMI 37.89 34.87 3.63 11.32 19.55 21.45 52.65 52.71 5.54 31.80 42.54 37.05
Vanilla FT 54.42 43.20 5.95 18.48 29.93 30.40 75.03 63.46 8.71 44.76 59.94 50.38
MTDNN 58.40 52.50 10.60 19.81 31.61 34.58 78.18 71.04 12.90 47.39 61.75 54.25
Ours 60.33 55.55 12.30 20.71 32.92 36.36 79.40 72.51 14.03 47.81 62.24 55.20
Table 2: Multiplex representation learning experiments on e-commerce networks: Clothes, Home, and Sports. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively.
Clothes Home Sports
Model cop cov bt cob Avg. cop cov bt cob Avg. cop cov bt cob Avg.
MPNet-v2 55.89 60.92 59.75 39.12 53.92 52.02 61.83 62.04 38.10 53.50 41.60 64.61 49.82 40.61 49.16
OpenAI-ada-002 65.30 70.87 69.44 48.32 63.48 60.99 71.43 71.36 47.86 62.91 50.80 73.70 60.20 54.06 59.69
DMGI 56.10 52.96 58.46 30.88 49.60 48.27 52.74 57.90 48.81 51.93 41.37 46.27 41.24 31.92 40.20
HDMI 62.85 63.00 69.69 52.50 62.01 51.75 57.91 57.91 53.39 55.24 45.43 61.22 55.56 52.66 53.72
Vanilla FT 81.57 80.46 88.52 67.38 79.48 73.72 75.49 85.80 76.83 77.96 68.22 77.11 80.78 78.46 76.14
MTDNN 80.30 78.75 87.58 65.94 78.14 72.49 75.17 84.00 77.29 77.24 66.20 76.50 79.72 78.69 75.28
Ours 82.04 81.18 88.90 68.34 80.12 73.59 79.06 86.58 80.07 79.83 67.92 79.85 81.52 81.54 77.71
5.3 Multiplex Representation Learning
We test the quality of the generated embeddings of different models on multiplex relation prediction tasks on the network. Given a query node/text (e.g., a paper or an item), a target relation (e.g., cb, sa, cop, or cov), and a candidate node/text list (e.g., papers or items), we aim to predict which key node in the candidate list should be linked to the given query node under the target relation. We use PREC@1 as the metric. More detailed information on experimental settings can be found in Appendix A.5.1
The results on academic networks and e-commerce networks are shown in Table 1 and Table 2, respectively. From the result, we can find that: 1) Metern performs significantly and consistently better than all the baseline methods on all types of relation prediction on all datasets (except cop prediction on Home and Sports). 2) In academic networks, multiplex representation learning methods (MTDNN and Metern) can always outperform the single representation learning method (Vanilla FT); while in e-commerce networks, Vanilla FT performs on par with MTDNN and Metern. This is because the relations in e-commerce networks are more semantically close to each other, while the relations in academic networks are less close to or even conflict with each other (see Figure 2 and Figure 7). 3) Although SPECTER Cohan et al. (2020), SciNCL Ostendorff et al. (2022), Sentence-Transformer Reimers and Gurevych (2019), and OpenAI-ada-002 Brown et al. (2020) are finetuned on large-scale corpora, they perform badly compared with Metern since they only output one embedding which cannot capture the diverse relation semantics. 4) Multiplex GNN methods (DMGI and HDMI) can utilize neighbor information from the network to output multiplex embeddings but fail to capture the diverse and dynamic text semantics associated with nodes.
5.4 Direct Inference with an Evident Source Relation
We conduct experiments on venue recommendation (Venue-Rec) and author identification (Author-Idf) in academic networks and brand prediction (Brand-Pred) in e-commerce networks to verify the effectiveness of Metern’s direct inference ability with an evident source relation. All three tasks are formulated as matching tasks where we use PREC@1 as the metric. For Venue-Rec, Author-Idf, and Brand-Pred, we use the sv, sa, and cob representation, respectively.
(a) Citation
(b) Rec
Figure 3: The learned source relation weights for citation prediction and paper recommendation on Geology network.
The results are shown in Table 3. From the results, we can find that: Metern outperforms all the baseline methods significantly and consistently on all the tasks (except Venue-Rec on Mathematics), which demonstrates that the multiplex embeddings generated by Metern are of good quality and can be directly applied to downstream tasks. Detailed information about the tasks and experimental settings can be found in Appendix A.5.2.
Table 3: Direct inference with an evident source relation (no task-specific training) on academic networks and e-commerce networks.
Geology Mathematics Clothes Home Sports
Model Venue-Rec Author-Idf Venue-Rec Author-Idf Brand-Pred Brand-Pred Brand-Pred
SPECTER 4.06 18.28 4.54 23.48 - - -
SciNCL 5.14 24.58 6.51 35.49 - - -
MPNet-v2 6.96 46.99 7.1 47.51 43.75 57.41 61.22
OpenAI-ada-002 6.20 45.35 6.07 42.74 59.09 63.59 69.14
DMGI 9.21 40.91 9.57 49.37 49.60 41.71 40.63
HDMI 2.96 20.36 3.44 25.55 53.07 58.30 51.07
Vanilla FT 11.72 62.76 13.35 68.31 66.16 74.36 76.87
MTDNN 13.63 63.72 15.03 68.45 63.79 77.98 78.86
Ours 14.44 68.43 14.64 71.63 66.69 79.50 81.87
5.5 Learning to Select Source Relations
We perform experiments on paper recommendation, paper classification, citation prediction, and year prediction on academic networks, and item classification and price prediction on e-commerce networks to test if Metern can learn to select source relations. We use PREC@1 as the metric for paper recommendation, Macro-F1 as the metric for paper classification and item classification, and RMSE as the metric for citation prediction, year prediction, and price prediction.
Table 4: Learn to select source relations on academic networks. ($\uparrow$) means the greater the score is, the better the model is, and ($\downarrow$) is on the contrary.
Geology Mathematics
Model Paper-Rec ($\uparrow$) Paper-Cla ($\uparrow$) Citation-Pred ($\downarrow$) Year-Pred ($\downarrow$) Paper-Rec ($\uparrow$) Paper-Cla ($\uparrow$) Citation-Pred ($\downarrow$) Year-Pred ($\downarrow$)
Vanilla FT $76.45_{0.00}$ $41.62_{0.04}$ $16.28_{0.01}$ $9.01_{0.01}$ $84.94_{0.00}$ $36.76_{0.18}$ $9.16_{0.01}$ $9.77_{0.00}$
MTDNN $77.99_{0.02}$ $45.84_{0.48}$ $\textbf{15.76}_{0.04}$ $8.63_{0.12}$ $85.42_{0.06}$ $38.47_{0.19}$ $9.13_{0.28}$ $9.74_{0.00}$
Ours $\textbf{80.58}_{0.00}$ $\textbf{46.68}_{0.08}$ $16.00_{0.09}$ $\textbf{8.34}_{0.01}$ $\textbf{86.56}_{0.03}$ $\textbf{39.20}_{0.29}$ $\textbf{8.95}_{0.02}$ $\textbf{9.26}_{0.03}$
Table 5: Learn to select source relations on e-commerce networks.
Clothes Home Sports
Model Item-Cla ($\uparrow$) Price-Pred ($\downarrow$) Item-Cla ($\uparrow$) Price-Pred ($\downarrow$) Item-Cla ($\uparrow$) Price-Pred ($\downarrow$)
Vanilla FT $87.84_{0.08}$ $17.88_{0.02}$ $90.68_{0.15}$ $76.41_{0.03}$ $\textbf{80.94}_{0.34}$ $20.27_{0.00}$
MTDNN $86.34_{0.38}$ $18.53_{0.16}$ $90.62_{0.20}$ $75.07_{0.06}$ $78.76_{0.14}$ $21.28_{0.05}$
Ours $\textbf{88.19}_{0.22}$ $\textbf{17.73}_{0.01}$ $\textbf{92.25}_{0.15}$ $\textbf{74.55}_{0.05}$ $80.37_{0.22}$ $\textbf{20.01}_{0.02}$
The results on academic networks and e-commerce networks are shown in Table 4 and Table 5, respectively. We also show the learned source relation weight for citation prediction and paper recommendation in Figure 3. From the results, we can find that: 1) Metern outperforms all the baseline methods significantly and consistently on all the tasks (except citation prediction on Geology and item classification on Sports). 2) Metern can learn to select different relations for different downstream tasks, i.e.,same-author” and “same-venue” for citation prediction (author and venue can largely determine the citation of the paper), and “cited-by” and “co-cited-by” for paper recommendation (paper in “cited-by” or “co-cited-by” relations are more likely to be beneficial to researchers together). Detailed information about the tasks and experimental settings can be found in Appendix A.5.3. More results analysis can be found in Appendix A.7.
5.6 Efficiency Analysis
Table 6: Time and memory costs on Geology.
Attribute Vanilla FT MTDNN Metern
Time 15h 35min 16h 23min 17h 20min
Memory 24,433 MB 26,201MB 28,357 MB
We run experiments on Geology to study the time complexity and memory complexity of Metern, comparing with Vanilla FT and MTDNN. All compared models are trained for 40 epochs on four Nvidia A6000 GPU devices with a total training batch size set as 512. We show the result in Table 6. From the result, we can find that: the time complexity and memory complexity of training Metern is on par with training Vanilla FT and MTDNN.
5.7 Multiplex Embedding Visualization
We visualize the multiplex node embeddings ${\bm{h}}_{v|r}$ learned by Metern with t-SNE Van der Maaten and Hinton (2008). The results on Geology network are shown in Figure 4. We randomly select one center node from the network, together with its neighboring nodes of the “cited-by” (cb), “same-author” (sa), and “same-venue” (sv) relations. Then, we utilize Metern to encode all the nodes (including the center node and all neighboring nodes) in different relation representation spaces (i.e., to obtain ${\bm{h}}_{v|r_{cb}}$, ${\bm{h}}_{v|r_{sa}}$, and ${\bm{h}}_{v|r_{sv}}$). In the figure, neighboring nodes of different relations are marked in different colors and shapes. From Figure 4, we can find that, in the output embedding space corresponding to one relation (e.g., the ${\bm{h}}_{v|r_{cb}}$ latent space corresponding to “cited-by” in Figure 4(a)), the center node (the blue circle)’s neighboring nodes of the corresponding relation (e.g., “cited-by” relation neighbors, marked as orange rectangles) are close to the center node. This demonstrates that Metern can capture the semantics of different relations and represent them in different latent subspaces.
(a) ${\bm{h}}_{v|r_{cb}}$
(b) ${\bm{h}}_{v|r_{sa}}$
(c) ${\bm{h}}_{v|r_{sv}}$
Figure 4: Multiplex embedding visualization on Geology network. cb, sa, and sv represent the “cited-by”, “same-author”, and “same-venue” relation respectively.
5.8 Analysis of Relation Weight $w_{r}$
In Metern, different relations sometimes have different learning speeds, and $w_{r}$ in Eq.(6) can be tuned to balance the learning speed of different relations. We analyze the effect of $w_{r}$ by showing the embedding performance on the validation set of two different weight settings (uniform and tuned) on the Geology network. The results are shown in Figure 5. In the uniform setting, all relations have the same weight (i.e., $[w_{cb},w_{sa},w_{sv},w_{cr},w_{ccb}]=[1,1,1,1,1]$). We can find that as the learning step increases, cb, cr, and ccb embeddings quickly reach the best performance and then overfit, while sa and sv embeddings are still underfitting. This motivates us to increase the weight for sa and sv since their learning speeds are slow. We can find that when we use the tuned weight set ($[w_{cb},w_{sa},w_{sv},w_{cr},w_{ccb}]=[1,2,2,1,1]$), the learning speeds of different relations are more balanced, leading to a better general multiplex encoder.
(a) cb
(b) sa
(c) sv
(d) cr
(e) ccb
Figure 5: Analysis of relation weights $w_{r}$ on Geology. The x-axis is the training step and the y-axis is MRR on the validation set. We compare uniform $w_{r}$ with a better-tuned $w_{r}$.
6 Related Work
6.1 Learning Embeddings on Multiplex Networks
Multiplex networks are also referred to as multi-view networks Qu et al. (2017); Shi et al. (2018) and multi-dimensional networks Berlingerio et al. (2013); Ma et al. (2018) in the literature. Multiplex networks consist of multiple relation/edge types among a set of single-type nodes. They can be seen as special cases of heterogeneous networks Dong et al. (2020); Sun et al. (2011) where nodes only have one type. The first category of existing works proposes to use one embedding vector to represent each node Dong et al. (2017); Hu et al. (2020); Wang et al. (2019); Zhang et al. (2019). HetGNN Zhang et al. (2019) introduces a graph neural network that utilizes heterogeneous neighbor aggregation to capture the diverse relation semantics. HAN Wang et al. (2019) proposes two-level attention including node-level and semantic-level to conduct relation semantics encoding. HGT Hu et al. (2020) introduces a more complex heterogeneous graph Transformer architecture to conduct node and edge semantics joint encoding. However, they all have an assumption that one embedding is enough to capture the semantics of all relations, which does not quite hold as shown in Section 3. The second category of works tries to develop multiplex embeddings for each node Jing et al. (2021); Park et al. (2020); Qu et al. (2017); Zhang et al. (2018), with one for each relation. MNE Zhang et al. (2018) proposes a random walk-based multiplex embedding learning method. MVE Qu et al. (2017) introduces a collaboration framework to capture the individual view network semantics and integrate them for the robust representations. DMGI Park et al. (2020) proposes a deep infomax framework to conduct multiplex embedding learning with graph neural networks. HDMI Jing et al. (2021) further enhances DMGI by introducing high-order mutual information with a fusion module. However, those works emphasize more on network structure encoding, while do not take the rich textual information associated with nodes into consideration.
6.2 Text Embeddings
Text embeddings Brown et al. (2020); Cohan et al. (2020); Harris (1954); Le and Mikolov (2014); Ostendorff et al. (2022); Reimers and Gurevych (2019) effectively capture the textual semantic similarity between text units (i.e., sentences and documents) via distributed representation learning. Early work Harris (1954) proposes the bag-of-words vector space model, representing a text unit as a multiset of its words, disregarding grammar and even word order. Paragraph Vector Le and Mikolov (2014) is then introduced to capture the text semantics rather than just word appearance by representing each document with a dense vector trained to predict words in the document. As large-scale pretrained language models Beltagy et al. (2019); Brown et al. (2020); Devlin et al. (2019); Liu et al. (2019b) are proposed, Sentence-BERT Reimers and Gurevych (2019) further finetunes BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019b) by using a siamese and triplet network structure to derive semantically meaningful text embeddings which can measure text proximity with the cosine similarity between embeddings. OpenAI-ada-002 is recently developed as a powerful text embedding model based on the GPT series of large language models Brown et al. (2020). Specific to the academic domain, SPECTER Cohan et al. (2020) finetunes SciBERT by introducing positive and negative paper pairs, while SciNCL Ostendorff et al. (2022) further enhances SPECTER by developing a nearest neighbor sampling strategy. However, all the existing works presume that one embedding can capture the general semantics for each text unit and does not take the diverse text relation semantics into consideration.
7 Conclusions
We discuss the limitations of our work in Appendix A.1. In this work, we tackle the problem of representation learning on multiplex text-rich networks. To this end, we introduce Metern, learning multiplex text representations with only one text encoder. Metern introduces relation prior tokens to capture the relation-specific signals and one text encoder to model the shared knowledge across relations. We conduct experiments on nine downstream tasks and five networks from two domains, where Metern outperforms baselines significantly and consistently. Interesting future directions include: (1) exploring other more advanced graph-empowered text encoders for learning multiplex embeddings, and (2) applying the framework to more network-related tasks such as network generation conditioned on various relations.
Appendix A Appendix
A.1 Limitations
In this work, we mainly focus on learning multiplex text/node representations on text-rich networks and solving downstream tasks (e.g., classification, regression, matching) with the learned embeddings. Because of the limited computational budgets, our PLM text encoder (bert-base-uncased) is medium-scale. In the future, we will explore applying a similar multiplex representation learning philosophy to large-scale language models. Other interesting future directions include designing more advanced graph-empowered language models to learn the multiplex embeddings and adopting the model into more real-world applications such as network generation.
A.2 Ethical Considerations
PLMs have been shown to be highly effective in encoding contextualized semantics and understanding documents, as evidenced by several studies Devlin et al. [2019], Liu et al. [2019b], Clark et al. [2020]. However, some researchers have pointed out certain limitations associated with these models, including the presence of social bias Liang et al. [2021] and the propagation of misinformation Abid et al. [2021]. In our work, we focus on utilizing the relation signals between texts from the multiplex text-rich network structure to facilitate the understanding of the semantics of the texts, which we believe could help to address issues related to bias and misinformation.
A.3 Datasets
The statistics of the five datasets can be found in Table 7. In academic networks, nodes correspond to papers and there are five types of relations between papers: “cited-by” (cb), “same-author” (sa), “same-venue” (sv), “co-reference” (cr), and “co-cited-by” (ccb); while in e-commerce networks, nodes are items and there are four types of relations between items: “co-purchased” (cop), “co-viewed” (cov), “bought-together” (bt), and “co-brand” (cob).
Table 7: Dataset Statistics.
Dataset #Nodes #Relations (Edges)
Geology 431,834 cb (1,000,000), sa (1,000,000), sv (1,000,000)
cr (1,000,000), ccb (1,000,000)
Mathematics 490,551 cb (1,000,000), sa (1,000,000), sv (1,000,000)
cr (1,000,000), ccb (1,000,000)
Clothes 208,318 cop (100,000), cov (100,000)
bt (100,000), cob (50,000)
Home 192,150 cop (100,000), cov (100,000)
bt (50,000), cob (100,000)
Sports 189,526 cop (100,000), cov (100,000)
bt (50,000), cob (100,000)
A.4 Distribution Shift between Different Relations
In Section 3, we show the learned embedding distribution shift between different relations on Geology in Figure 2. In this section, we calculate the raw data distribution shift between different relations’ distribution $P_{r_{k}}(e_{ij}|v_{i},v_{j})$. The distribution shift is measured by Jaccard score:
$\text{Jac}(r_{k},r_{l})=\frac{|P_{r_{k}}(e_{ij})\cap P_{r_{l}}(e_{ij})|}{|P_{r_{k}}(e_{ij})\cup P_{r_{l}}(e_{ij})|}$ (11)
Since the whole networks are too large to calculate the Jaccard score, we randomly sample a sub-network from each network that contains 10,000 nodes and calculate Eq.(11). The results on Geology, Mathematics, Clothes, Home and Sports networks can be found in Figure 6. If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly one, which is not the case in Figure 6.
More empirical experiments on the learned embedding distribution shift between relations in Mathematics, Clothes, Home, and Sports networks can be found in Figure 7. We finetune BERT [^4] We use the bert-base-uncased checkpoint. Devlin et al. [2019] to generate embeddings on one source relation distribution $P_{r_{k}}(e_{ij}|v_{i},v_{j})$ (row) and test the embeddings on the same or another target relation distribution $P_{r_{l}}(e_{ij}|v_{i},v_{j})$ (column). If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly the same, which is not the case in Figure 7.
(a) Geology
(b) Mathematics
(c) Clothes
(d) Home
(e) Sports
Figure 6: Raw data distribution shift between different relations on Geology, Mathematic, Clothes, Home, and Sports network. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively. Each entry is the Jaccard score between the corresponding two relation distributions.
(a) Mathematics
(b) Clothes
(c) Home
(d) Sports
Figure 7: Distribution shift between different relations on Mathematic, Clothes, Home, and Sports network. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively. Each entry is the PREC@1 of BERT fine-tuned on the corresponding source relation distribution and tested on the corresponding target relation distribution.
A.5 Experimental Setting
A.5.1 Multiplex Representation Learning
Hyperparameter setting.
To facilitate the reproduction of our representation learning experiments, we provide the hyperparameter configuration in Table 8. Vanilla FT, MTDNN, and Metern use exactly the same set of hyperparameters for a fair comparison. The last layer [CLS] token hidden states are utilized to develop ${\bm{h}}_{v|r}$ for Vanilla FT, MTDNN, and Metern. Paper titles and item titles are used as text associated with the nodes in the two kinds of networks, respectively. (For some items, we concatenate the item title and description together since the title is too short.) The models are trained for 40 epochs on 4 Nvidia A6000 GPUs with a total batch size of 512. The total time cost is around 17 hours and 2 hours for networks in the academic domain and e-commerce domain respectively. Code is available at https://anonymous.4open.science/r/METER-submit-2C7B.
Table 8: Hyper-parameter configuration for representation learning.
Parameter Geology Mathematics Clothes Home Sports
Max Epochs 40 40 40 40 40
Peak Learning Rate 5e-5 5e-5 5e-5 5e-5 5e-5
Batch Size 512 512 512 512 512
# Prior Tokens $m$ 5 5 5 5 5
 Warm-Up Epochs 4 4 4 4 4
Sequence Length 32 32 32 32 32
Adam $\epsilon$ 1e-8 1e-8 1e-8 1e-8 1e-8
Adam $(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999) (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
Clip Norm 1.0 1.0 1.0 1.0 1.0
Dropout 0.1 0.1 0.1 0.1 0.1
A.5.2 Direct Inference with an Evident Source Relation
We provide problem definitions and experimental settings for tasks in Section 5.4. The tasks include venue recommendation, author identification, and brand prediction.
Venue Recommendation.
Problem Definition. Given a query paper node (with associated text) and a candidate venue list (each with its published papers), we aim to predict which venue in the candidate list should be recommended for the given query paper.
Experimental Settings. We adopt in-batch testing with a testing batch size of 256. We use PREC@1 as the metric. The max sequence length is 32 and 256 for the query paper and venue (concatenation of its 100 randomly selected published papers’ titles) respectively.
Author Identification.
Problem Definition. Given a query paper node (with associated text) and a candidate author list (each with his/her published papers), we aim to predict which people in the candidate list is the author of the given query paper.
Experimental Settings. We adopt in-batch testing with a testing batch size of 256. We use PREC@1 as the metric. The max sequence length is 32 and 256 for the query paper and author (concatenation of his/her 100 randomly selected published papers’ titles) respectively.
Brand Prediction.
Problem Definition. Given a query item node (with associated text) and a candidate brand list (each with items in the brand), we aim to predict which one in the candidate list is the brand for the given query item.
Experimental Settings. We adopt in-batch testing with a testing batch size of 256. We use PREC@1 as the metric. The max sequence length is 32 and 256 for query paper and brand (concatenation of its 100 randomly selected items’ texts) respectively.
A.5.3 Learn to Select Source Relations
We provide problem definitions and experimental settings for tasks in Section 5.5. The tasks include paper recommendation, paper classification, citation prediction, year prediction, item classification, and price prediction.
Paper Recommendation.
Problem Definition. Given a query paper node (with associated text) and a candidate paper list (each with associated text), we aim to predict which paper in the candidate list should be recommended to the people who are interested in the query paper.
Experimental Settings. We have 1,000 samples in the train set to teach the models how to select source relations, 1,000 samples in the validation set to conduct the early stop, and 100,000 samples in the test set to evaluate the performance of the models. The learning rate is set as 1e-3, the training batch size is 128, and the testing batch size is 256. We conduct the in-batch evaluation with PREC@1 as the metric. All experiments are done on one NVIDIA A6000. We repeat three runs for each model and show the mean and standard deviation in Table 4.
Paper Classification.
Problem Definition. Given a query paper node (with associated text), we aim to predict what is the category of the paper. The number of paper node categories in academic networks (Geology and Mathematics) is shown in Table 9.
Table 9: Number of paper node categories in academic networks.
Geology Mathematics
18 17
Experimental Settings. We have 1,000 samples for each category in the train set to teach the models how to select source relations, 200 samples for each category in the validation set to conduct the early stop, and 200 samples for each category in the test set to evaluate the performance of the models. The learning rate is tuned in 5e-3 and 1e-3, the training batch size is 256, and the testing batch size is 256. We adopt Macro-F1 as the metric. All experiments are done on one NVIDIA A6000. We repeat three runs for each model and show the mean and standard deviation in Table 4.
Citation Prediction.
Problem Definition. Given a query paper node (with associated text), we aim to predict its future number of citations.
Experimental Settings. For both Geology and Mathematics datasets, we extract papers the citation of which is in the range from 0 to 100. We randomly select 10,000 papers from the extracted papers to form the training set, 2,000 papers to form the validation set, and 2,000 papers to form the test set. The learning rate is set as 1e-2 for all compared methods, the training batch size is 256, and the testing batch size is 256. We adopt RMSE as the metric. All experiments are done on one NVIDIA A6000. We repeat three runs for each model and show the mean and standard deviation in Table 4.
Year Prediction.
Problem Definition. Given a query paper node (with associated text), we aim to predict the year when it was published.
Experimental Settings. For both Geology and Mathematics datasets, we conduct minus operations to make the smallest ground truth year to be 0 (we minus all year numbers by the earliest year, i.e., 1981, in MAG). We randomly select 10,000 papers from the extracted papers to form the training set, 2,000 papers to form the validation set, and 2,000 papers to form the test set. The learning rate is set as 1e-2 for all compared methods, the training batch size is 256, and the testing batch size is 256. We adopt RMSE as the metric. All experiments are done on one NVIDIA A6000. We repeat three runs for each model and show the mean and standard deviation in Table 4.
Item Classification.
Problem Definition. Given a query item node (with associated text), we aim to predict what is the category of the item. The number of item node categories in e-commerce networks (Clothes, Home, and Sports) is shown in Table 10.
Table 10: Number of item node categories in e-commerce networks.
Clothes Home Sports
7 9 16
Experimental Settings. We have 1,000 samples for each category in the train set to teach the models how to select source relations, 200 samples for each category in the validation set to conduct the early stop, and 200 samples for each category in the test set to evaluate the performance of the models. The learning rate is tuned in 5e-3 and 1e-3, the training batch size is 256, and the testing batch size is 256. We adopt Macro-F1 as the metric. All experiments are done on one NVIDIA A6000. We repeat three runs for each model and show the mean and standard deviation in Table 5.
Price Prediction.
Problem Definition. Given a query item node (with associated text), we aim to predict its price.
Experimental Settings. For Clothes, Home, and Sports, we delete the long-tail items and keep items whose prices are under 100/1,000/100 respectively. We randomly select 10,000 items from the extracted items to form the training set, 2,000 items to form the validation set, and 2,000 items to form the test set. The learning rate is set as 1e-2 for all compared methods, the training batch size is 256, and the testing batch size is 256. We adopt RMSE as the metric. All experiments are done on one NVIDIA A6000. We repeat three runs for each model and show the mean and standard deviation in Table 5.
A.6 Relation Embedding Initialization Study.
We study how different relation embedding initialization affects the quality of multiplex representations learned by Metern. We explore three initialization settings: zero vectors initialization, normal distribution initialization, and word embedding initialization. The results of average PREC@1 on different networks are shown in Table 11. From the results, there is no significant difference between the representation learning quality of different initialized relation embeddings.
Table 11: Performance of different relation embedding initialization on different networks.
Model Geology Mathematics Clothes Home Sports
Metern w/ zero init 36.31 55.26 79.99 79.68 77.59
Metern w/ randn init 36.43 55.20 80.19 79.69 77.56
Metern w/ word init 36.36 55.20 80.12 79.83 77.71
A.7 More Results on Learn to Select Source Relations.
In section 4.2, we propose to let the model learn to select source relations for different downstream tasks and show the learned source relation weight for citation prediction and paper recommendation in Figure 3. In this section, we show more results on the learned source relation weight on academic network downstream tasks in Figure 8 and on e-commerce network downstream tasks in Figure 9. There are four downstream tasks in the academic domain: year prediction, citation prediction, paper classification, and paper recommendation. There are two downstream tasks in the e-commerce domain: price prediction and item classification. From Figure 8 and Figure 9, we can find that different relations can benefit different downstream tasks.
On the year prediction task, the “co-cited-by” relation and “same-author” relation are more useful. This indicates that papers tend to cite recent papers and authors tend to be active within a short period (e.g., active during the Ph.D. study and stop publishing papers after graduation).
On the citation prediction task, the “same-author” relation and “same-venue” relation are more beneficial. This implies that the impact of papers is more determined by the published venue and the author who writes them (people tend to follow works from famous researchers).
On the paper classification task, the “cited-by” relation is quite useful. This means that papers and their cited papers have a tendency to have the same fine-grained topics.
On the paper recommendation task, the “cited-by” relation and “co-cited-by” relation dominate. The goal of the paper recommendation task is to recommend similar papers to researchers which may contribute to their own research development. The result is interesting since papers and their cited papers have a tendency to have the same topic, and thus should be recommended to researchers together, and papers in the “co-cited-by” relation have already demonstrated that their ideas can be combined and result in a new paper (they are cited by same papers), and thus should be recommended together.
On the price prediction task, the “co-viewed” and “bought-together” relation matters a lot. This implies that the same user tends to view items and buy together items of a similar price range.
On the item classification task, the “co-viewed” relation dominates. This means that items co-viewed by the same user tend to have similar functions.
(a) Geology: Year
(b) Geology: Citation
(c) Geology: Class
(d) Geology: Rec
(e) Mathematics: Year
(f) Mathematics: Citation
(g) Mathematics: Class
(h) Mathematics: Rec
Figure 8: Learnt source relation weights for tasks on academic networks (Geology and Mathematics). The x-axis is the learned weight vector id and the y-axis is the relation embedding id grouped by relation id. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively.
(a) Clothes: Price
(b) Clothes: Class
(c) Home: Price
(d) Home: Class
(e) Sports: Price
(f) Sports: Class
Figure 9: Learnt source relation weights for tasks on e-commerce networks (Clothes, Home, and Sports). The x-axis is the learned weight vector id and the y-axis is the relation embedding id grouped by relation id. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively.
A.8 Relation Token Embedding Visualization.
We visualize the relation token embeddings $Z_{\mathcal{R}}$ learned by Metern with t-SNE Van der Maaten and Hinton [2008], projecting those embeddings into a 2-dimensional space. The results on the Geology network and Mathematics network are shown in Figure 10, where the embeddings belonging to the same relation are assigned the same color. From the results, we can find that embeddings belonging to the same relation are close to each other, while those belonging to different relations are discriminative. This demonstrates that Metern can learn to capture the different semantics of different relations by assigning different relations’ embeddings to different areas in the latent space.
(a) Geology
(b) Mathematics
Figure 10: Visualization of relation embedding $Z_{R}$ on Geology and Mathematics. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively. We can find that embeddings belonging to the same relation are close to each other, while those belonging to different relations are discriminative.